### 27 Dec 2023 v2

@@ Abstract
Advancements in clinical treatment are increasingly constrained by the limitations of supervised learning techniques, which depend heavily on large volumes of annotated data. The annotation process is not only costly but also demands substantial time from clinical specialists. Addressing this issue, we introduce the S4MI (Self-Supervision and Semi-Supervision for Medical Imaging) pipeline, a novel approach that leverages the advancements in self-supervised and semi-supervised learning. These techniques engage in auxiliary tasks that do not require labeling, thus simplifying the scaling of machine supervision compared to fully-supervised methods. Our study benchmarks these techniques on three distinct medical imaging datasets to evaluate their effectiveness in classification and segmentation tasks. Remarkably, we observed that self-supervised learning with only 10% of the annotation surpassed the performance of full annotation in the classification of most datasets. Similarly, the semi-supervised approach demonstrated superior outcomes in segmentation, outperforming fully-supervised methods with 50% fewer labels across all datasets. In line with our commitment to contributing to the scientific community, we have made the S4MI code openly accessible, allowing for broader application and further development of these methods. The code can be accessed at link.
@@ 3.1 Classification
We present the results of supervised and self-supervised trained classifiers in Table 1. We use the F1 score as the comparison metric, similar to its implementation in previous works [6]. The F1 score is defined as $F1=\frac{2*Precision*Recall}{Precision+Recall}=\frac{2*TP}{2*TP+FP+FN}$, where $TP$ represents True Positive, $FN$ is False Negative and $FP$ is False Positive. We observe that CASS (Figure 1) outperforms DINO and transfer learning consistently across all the datasets and backbones, with the, exception for the ISIC 2017 challenge dataset, where DINO outperforms CASS using the ResNet50 backbone. Interestingly, we also observe that ViT trained with CASS using just 10% labels, which saves significant annotation time, performs better than the transfer learning-based supervised approach with 100% labels. For the same case, DINO performs on par with the supervised approach while using 90% fewer labels. This shows the impact that machine supervision can have on improving access to classified medical images. DINO and CASS take an image to create asymmetry through augmentation (DINO) and architecture (CASS). These are then passed through different parameterized feature extractors (DINO) and architectural feature extractors (CASS) to create embeddings. Since the two embeddings are generated from the same image, they are expected to be similar; this is then used as the supervisory signal to maximize the similarity between the two embedding and update encoders. In CASS, architectural invariance is used to create a positive pair instead of creating augmentation invariance, as in the case of DINO. This is because CNN and Transformers learn different representations from the same image [6].
@@ 4 Discussion
We studied three training regimes for segmentation: supervised, semi-supervised, and unsupervised. While applying the unsupervised approach requires some adapta- tion and hyperparameter tuning, the dermofit data set consistently shows that semi-supervised outperforms fully-supervised, regardless of labeling percentage. However, we con- sistently improve performance using a semi-supervised approach – wherein we used labeled and unlabeled data per batch to help the model learn better representation. The result on the dermatomyositis dataset using no labels is close as compared to the supervised and semi-supervised with 10% labels. Between 50% and 70% labeling, the performance of semi-supervised and fully-supervised merges. However, the result on the ISIC-2017 data set with 10% labels indicates fully-supervised performs better than the semi-supervised; however as the percentage of labels increases above 10% semi-supervised outperforms fully-supervised. Additionally, unlike the unsupervised approach, the supervised and the semi-supervised approaches don’t treat this problem as pixel clustering and are less prone to overfitting to the dominant pixel distribution. Since classification requires image-level identification of classes instead of pixel-level identification, segmentation is a more difficult objective. Yet, introducing reduced human supervision with pseudo-labels in the semi-supervised approach improved performance over the supervised and the unsupervised paradigms to the where the semi-supervised approach was able to outperform the fully supervised approach with 50% fewer labels.

### 18 Mar 2024 v3

@@ Abstract
Advancements in clinical treatment are increasingly constrained by the limitations of supervised learning techniques, which depend heavily on large volumes of annotated data. The annotation process is not only costly but also demands substantial time from clinical specialists. Addressing this issue, we introduce the S4MI (Self-Supervision and Semi-Supervision for Medical Imaging) pipeline, a novel approach that leverages the advancements in self-supervised and semi-supervised learning. These techniques engage in auxiliary tasks that do not require labeling, thus simplifying the scaling of machine supervision compared to fully-supervised methods. Our study benchmarks these techniques on three distinct medical imaging datasets to evaluate their effectiveness in classification and segmentation tasks. Notably, we observed that self-supervised learning significantly surpassed the performance of supervised methods in the classification of all evaluated datasets. Remarkably, the semi-supervised approach demonstrated superior outcomes in segmentation, outperforming fully-supervised methods while using 50% fewer labels across all datasets. In line with our commitment to contributing to the scientific community, we have made the S4MI code openly accessible, allowing for broader application and further development of these methods. The code can be accessed at link.
@@ 3.1 Classification
We present the results of supervised and self-supervised trained classifiers in Table 1. We use the F1 score as the comparison metric, similar to its implementation in previous works [6]. The F1 score is defined as $F1=\frac{2*Precision*Recall}{Precision+Recall}=\frac{2*TP}{2*TP+FP+FN}$, where $TP$ represents True Positive, $FN$ is False Negative and $FP$ is False Positive. We observe that CASS (Figure 1) outperforms DINO and transfer learning consistently across all the datasets and backbones, with the, exception for the ISIC 2017 challenge dataset, where DINO outperforms CASS using the ResNet50 backbone. Interestingly, we also observe that ViT trained with CASS using just 10% labels, which saves significant annotation time, performs significantly better than the transfer learning-based supervised approach with 100% labels for the Dermatomyositis dataset. For the same case, DINO performs on par with the supervised approach while using 90% fewer labels. This shows the impact that machine supervision can have on improving access to classified medical images. DINO and CASS take an image to create asymmetry through augmentation (DINO) and architecture (CASS). These are then passed through different parameterized feature extractors (DINO) and architectural feature extractors (CASS) to create embeddings. Since the two embeddings are generated from the same image, they are expected to be similar; this is then used as the supervisory signal to maximize the similarity between the two embedding and update encoders. In CASS, architectural invariance is used to create a positive pair instead of creating augmentation invariance, as in the case of DINO. This is because CNN and Transformers learn different representations from the same image [6].
@@ 4 Discussion
We studied three training regimes for segmentation: supervised, semi-supervised, and unsupervised. While applying the unsupervised approach requires some adaptation and hyperparameter tuning, the dermofit data set consistently shows that semi-supervised outperforms fully-supervised, regardless of labeling percentage. However, we consistently improve performance using a semi-supervised approach – wherein we used labeled and unlabeled data per batch to help the model learn better representation. The result on the dermatomyositis dataset using no labels is close as compared to the supervised and semi-supervised with 10% labels. Between 50% and 70% labeling, the performance of semi-supervised and fully-supervised merges. However, the result on the ISIC-2017 data set with 10% labels indicates fully-supervised performs better than the semi-supervised; however as the percentage of labels increases above 10% semi-supervised outperforms fully-supervised. Additionally, unlike the unsupervised approach, the supervised and the semi-supervised approaches don’t treat this problem as pixel clustering and are less prone to overfitting to the dominant pixel distribution. Since classification requires image-level identification of classes instead of pixel-level identification, segmentation is a more difficult objective. Yet, introducing reduced human supervision with pseudo-labels in the semi-supervised approach improved performance over the supervised and the unsupervised paradigms to the where the semi-supervised approach was able to outperform the fully supervised approach with 50% fewer labels.